# Motility-Driven Viscoelastic Control of Tissue Morphology in Presomitic Mesoderm

### v1

@@ VII Acknowledgements
Supplementary Figure S4: (a) Standard Linear Fluid II model representing viscoelastic behavior, consisting of a dashpot with viscosity $\eta_{1}$ in series with a Kelvin–Voigt element (spring with modulus $E$ and dashpot with viscosity $\eta_{2}$). (b) Time profile of the applied mechanical perturbation.
We investigated the response of a standard viscoelastic fluid subjected to spatially pulsatile perturbations, using the Standard Linear Fluid II model Kelly (2025). In this framework, a dashpot (viscosity $\eta_{1}$) is connected in series with a Kelvin–Voigt element, comprising a spring (modulus $E$) and a dashpot (viscosity $\eta_{2}$)(Fig S4,(a)). The resulting stress–strain relationship is given by:
@@
Where:
  • $\tilde{\boldsymbol{\sigma}}$ is the stress tensor,
  • $\tilde{\boldsymbol{\varepsilon}}=\frac{1}{2}\left(\nabla\mathbf{u}+\nabla\mathbf{u}^{\top}\right)$ is the linearized strain tensor,
  • $\mathbf{u}(\mathbf{x},t)$ is the displacement vector field,
  • $E$ is the elastic modulus,
  • $\eta_{1}$ and $\eta_{2}$ are the viscosity. (Fig. S4(a)).
@@
$\Theta(t)=\sum_{n=0}^{\infty}\text{rect}\left(\frac{t-n(\mathrm{T_{on}}+\mathrm{T_{off}})-\frac{\mathrm{T_{on}}}{2}}{\mathrm{T_{on}}}\right)$ (S16)
@@
Since, this residual deformation is going to accumulate in every cycle without affecting any scaling behavior. So we can write the total deformation at the end of cycle $n$,
$\hat{u}_{n\gg 1}(k_{0},\tau_{1})\sim\frac{n}{k_{0}^{2}\tau_{1}}$ (S34)
To establish equivalence with our simulations, we identify the stress relaxation timescale $\tau_{s}$ from simulations with the viscous relaxation time $\tau_{1}$ in the Standard Fluid model. This identification is supported by the matched temporal decay of stress in both simulation data and analytical solutions, confirming that $\tau_{s}$ and $\tau_{1}$ characterize the same relaxation process. Based on this, the defomation becomes,
$\hat{u}_{n\gg 1}(k_{0},\tau_{1})\sim\frac{1}{k_{0}^{2}\tau_{s}}$ (S35)

### v2

@@ VII Acknowledgements
Figure S4: (a) Linear viscoelastic model with two timescales, consisting of a dashpot with viscosity $\eta_{1}$ in series with a Kelvin–Voigt element (spring with modulus $E$ and dashpot with viscosity $\eta_{2}$). (b) Time profile of the applied mechanical perturbation.
We investigated the response of a standard viscoelastic fluid subjected to spatially pulsatile perturbations, using a linear model with two timescales kellyMechanicsLectureNotes2025. In this framework, a dashpot (viscosity $\eta_{1}$) is connected in series with a Kelvin–Voigt element, comprising a spring (modulus $E$) and a dashpot (viscosity $\eta_{2}$) in parallel(Fig S4,(a)). The resulting stress–strain relationship is given by:
@@
Where $\tilde{\boldsymbol{\sigma}}$ is the stress tensor, $\tilde{\boldsymbol{\varepsilon}}=\frac{1}{2}\left(\nabla\mathbf{u}+\nabla\mathbf{u}^{\top}\right)$ is the linearized strain tensor, $\mathbf{u}(\mathbf{x},t)$ is the displacement vector field, $E$ is the elastic modulus, $\eta_{1}$ and $\eta_{2}$ are the viscosity. (Fig. S4(a)).
@@
$\Theta(t)=\sum_{n=0}^{\infty}\operatorname{rect}\!\left(\frac{t-n(\mathrm{T_{on}}+\mathrm{T_{off}})-\frac{\mathrm{T_{on}}}{2}}{\mathrm{T_{off}}}\right)$ (S16)
$\operatorname{rect}(x)=\begin{cases}1,&|x|\leq\tfrac{1}{2},\\[6.0pt] 0,&|x|>\tfrac{1}{2}.\end{cases}$
@@
Since this residual deformation accumulates in every cycle without affecting any scaling behavior, we can write the total deformation at the end of cycle $n$,
$\hat{u}_{n}(k_{0},\tau_{1})\sim\frac{n}{k_{0}^{2}\tau_{1}}$ (S34)
To establish equivalence with our simulations, we identify the stress relaxation timescale $\tau_{s}$ from simulations with the viscous relaxation time $\tau_{1}$ in the linear viscoelastic model. Hence,
$\hat{u}_{n}(k_{0},\tau_{1})\sim\frac{1}{k_{0}^{2}\tau_{s}}$ (S35)